A spring balance reads 10 kg in a lift when the lift is stationary. When the lift starts moving with a
constant acceleration, the new reading is 12.3 kg. If the upward acceleration is considered positive,
what is the acceleration of the lift? Acceleration due to gravity may be taken as 10
m/s2downwards

Grade:

1 Answers

Aziz Alam IIT Roorkee
9 years ago
\text{N=mg ,when there is no acceleration of system apart from gravity}\text{N'=m(g+a) , acceleration of the system is also included}
\text{Reading by weighing machine, M'=N'/g}
\text{a=g*(M'/M-1)} \Rightarrow a=2.3 m/s^2